### v2

# WinoPron: Revisiting English Winogender Schemas for Consistency, Coverage, and Grammatical Case

Abstract
While measuring bias and robustness in coreference resolution are important goals, such measurements are only as good as the tools we use to measure them with. Winogender schemas (Rudinger et al., 2018) are an influential dataset proposed to evaluate gender bias in coreference resolution, but a closer look reveals issues with the data that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, violations of template constraints, and typographical errors. We identify these issues and fix them, contributing a new dataset: WinoPron. Our changes affect performance with state-of-the-art supervised coreference resolution systems as well as all model sizes of the language model FLAN-T5, with F1 dropping on average 10 percentage points. We also propose a new method to evaluate pronominal bias in coreference resolution that goes beyond the binary. With this method and our new dataset which is balanced for grammatical case, we empirically demonstrate that bias characteristics vary not just across pronoun sets, but also across surface forms of those sets.
1 Introduction
Third-person pronouns (he, she, they, etc.) help us refer to people in conversation. Since they mark referential gender in English, gender bias affects how coreference resolution systems map these pronouns to people. Rudinger et al. (2018) demonstrated this by introducing Winogender schemas, a challenge dataset to evaluate occupational gender bias in coreference resolution systems. The dataset has become popular due to its careful construction; it has been translated to other languages (Hansson et al., 2021; Stanovsky et al., 2019) and used in framings beyond coreference resolution, e.g., to evaluate natural language inferences (Poliak et al., 2018), intrinsic bias in language models (Kurita et al., 2019), and pronoun fidelity (Gautam et al., 2024).
However, a closer look at the dataset reveals weaknesses that compromise its use for reliable evaluation, including treating different pronominal forms as equivalent, varying more than just pronouns in templates that should otherwise be identical, and not conforming to structural requirements.
In this paper, we identify issues with the original dataset and fix them to create a new, expanded WinoPron3). [^1] Data and code available at github.com/uds-lsv/winopron. We then empirically show how our fixes affect coreference resolution system performance (§4), and contribute a novel method to evaluate pronominal bias in coreference resolution (§5) that goes beyond the binary and focuses on linguistic gender without mapping it to social gender (Cao and Daumé III, 2021).
Figure 1: Performance on our new WinoPron dataset with supervised coreference resolution systems (CAW-coref and LingMess), and language models fine-tuned for coreference resolution (SpanBERT) and prompted zero-shot (FLAN-T5).
WinoPron is more difficult than Winogender schemas for state-of-the-art supervised coreference resolution systems as well as large language models untrained for the task. We find that grammatical case, which we balance for in WinoPron, matters for both performance and bias results; accusative pronouns are harder to resolve than nominative or possessive pronouns, and system pronominal bias is not always consistent across different grammatical cases of the same pronoun set.
2 Background: Winogender Schemas
[…]
(a) The cashier told the customer that his / her / their / xyr card was declined.
[…]
Figure 2: Winogender schemas for cashier, customer and possessive pronouns, with the antecedent bolded.
Winogender schemas (Rudinger et al., 2018) are a widely-used dataset consisting of paired sentence templates in English, with slots for two human entities (an occupation and a participant), and a third person singular pronoun. As Figure 2 shows, the second part of each template disambiguates which of the two entities the pronoun uniquely refers to. Changing the pronoun (e.g., from his to her) maintains the coreference, allowing us to measure whether coreference resolution systems are worse at resolving certain pronouns to certain entities. Rudinger et al. (2018) use the gendered associations of these pronouns to show that gender bias affects coreference resolution performance.
The entities consist of 60 occupation-participant pairs (e.g., accountant is paired with taxpayer; see Appendix B for a full list). A pair of templates is created for each occupation and participant, resulting in 120 templates. Each template can be instantiated with three pronoun sets (he, she, and singular they), for a total of 120 x 3 = 360 sentences.
3 WinoPron Dataset
Although Winogender schemas are established in the coreference resolution literature, we find issues in the dataset that compromise its use for reliable evaluation. We motivate them and describe our fixes here and in Appendix F, creating WinoPron. We reuse the occupation-participant pairs from Winogender schemas, but add 240 templates to cover missing grammatical cases for a total of 360 templates. We also include a neopronoun set (xe), giving us 360 templates x 4 pronouns = 1,440 sentences, which we use for evaluation in §4 and §5.
Grammatical case WS WP
Nominative (he, she, they, xe) 89 120
Accusative (him, her, them, xem) 4 120
Possessive (his, her, their, xyr) 27 120
Table 1: Number of templates per grammatical case in Winogender schemas (WS) and WinoPron (WP).
3.1 Issues and Solutions
Support for 3 Grammatical Cases
We hypothesize that systems have different performance and bias characteristics with pronouns in different grammatical cases. [^2] Here, we mean the surface form of the pronoun. However, as Table 1 shows, Winogender schemas have a variable number of pronouns per grammatical case, and treat them all as equivalent. To enable more granular evaluation, we balance this distribution in WinoPron.
Single-Entity Versions
When evaluating large language models on coreference resolution when they have not explicitly been trained for it, poor performance could mean that the model simply cannot perform the task (with a given prompt). In its current form, Winogender schemas do not allow us to disentangle why bad model performance is bad. In WinoPron, we create single-entity sentences that are parallel to the traditional, more complex double-entity sentences, for a simple setting to test this, and a useful baseline for all systems.
Dataset Data LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
WS All 0.85 0.81 0.72 0.82 0.52 0.67 0.89 0.97 0.97
WP All 0.64 0.67 0.62 0.70 0.52 0.62 0.78 0.89 0.89
Nominative 0.73 0.78 0.67 0.77 0.52 0.65 0.85 0.95 0.97
Accusative 0.52 0.57 0.55 0.60 0.50 0.58 0.70 0.83 0.79
Possessive 0.67 0.66 0.63 0.74 0.52 0.63 0.79 0.90 0.91
Grouped
PronounC 0.06 0.25 0.51 0.60 0.04 0.01 0.15 0.56 0.45
DisambigC 0.41 0.43 0.25 0.42 0.00 0.00 0.05 0.52 0.44
Table 2: F1 scores of coreference resolution systems on Winogender schemas (WS) and WinoPron (WP). We report total F1, performance split by grammatical case, and F1 consistency on templates grouped by pronouns (PronounC) and grouped by disambiguating pairs (DisambigC). Red, italicized numbers are worse than chance.
Support for All English Pronouns
For a controlled evaluation comparing pronouns, it is common to use templates that only vary the pronoun. However, 17% of Winogender schemas must be modified to work with singular they due to its different verbal agreement (“he was” but “they were”). To ensure a fair comparison between pronouns, we modify these templates to work with any pronouns, including neopronouns (Lauscher et al., 2022).
Consistency Fixes
Winograd-like schemas have strict structural constraints so that models cannot inflate performance through heuristics. However, when analyzing Winogender schemas, we found constraint violations, e.g., non-parallel paired templates. We fixed these along with typographical errors to ensure robust and reliable evaluation.
3.2 Validation
Two authors with linguistic training create and validate the sentence templates for grammaticality and correct, unique coreferences in an iterative process, i.e., we discussed, modified, and re-annotated templates as required until we reached consensus. An additional annotator independently verified the final templates, rating 100% of them as grammatical, and 98.2% as having unique, correct coreferences.
4 Performance and Consistency
To demonstrate the effects of our changes, we evaluate Winogender schemas and WinoPron with a range of models, including supervised coreference resolution systems and large language models.
4.1 Models
LingMess (Otmazgin et al., 2023) is a state-of-the-art, linguistically motivated, mixture-of-experts system for coreference resolution.
CAW-coref (D’Oosterlinck et al., 2023) is a state-of-the-art word-level coreference resolution system based on an encoder-only model.
SpanBERT (Joshi et al., 2020) is an encoder-only language model pre-trained with a span prediction objective and further enhanced for coreference resolution with fine-tuning data.
FLAN-T5 (Chung et al., 2024) is an instruction-tuned language model which is not trained for coreference resolution. We evaluate on five model sizes (small, base, large, xl, and xxl), with prompts from the FLAN collection (Longpre et al., 2023). See Appendix E for details on prompting.
Pronouns Nominative case Accusative case Possessive case
Positive Negative Positive Negative Positive Negative
he/him/his engineer painter receptionist secretary dietitian secretary practitioner chef hairdresser secretary
she/her/her hairdresser painter accountant plumber cashier firefighter mechanic practitioner painter accountant surgeon
they/them/their accountant plumber cashier b dietitian advisor baker accountant surgeon
xe/xem/xyr hairdresser engineer mechanic cashier advisor baker engineer supervisor
Table 3: A sample of SpanBERT-large’s biases when resolving pronouns to occupations. Positive bias: the model over-resolves the pronoun to that occupation. Negative bias: the model under-resolves the pronoun to the occupation.
4.2 Performance results
Table 2 shows how model performance changes with our new dataset. Firstly, WinoPron is consistently harder than Winogender schemas across all systems, with F1 dropping on average by 10 percentage points. Scaling behaviour generally holds for SpanBERT and FLAN-T5. In fact, scale seems to supercede supervision, as the largest FLAN-T5 models perform the best overall.
As we hypothesized, there are large performance differences across grammatical case: systems perform best at resolving nominative pronouns, with a slight decrease for possessive and a large drop for accusative pronouns. We report additional results on different pronoun sets in Appendix G.
4.3 Consistency results
We follow Ravichander et al. (2022) in evaluating system robustness via the consistency of outputs across groups of closely related instances, i.e., the score of the lowest-performing instance in the group is taken as the group’s score. We consider two groups: (a) pronoun consistency, and (b) disambiguation consistency. Pronoun consistency measures whether a template can be solved regardless of the pronoun it is filled with. Since we consider four pronoun sets, chance is $0.5^{4}$, or $0.0625$. Inspired by pair accuracy on Winograd schemas  (Abdou et al., 2020), we also measure disambiguation consistency on WinoPron. Here, we keep the pronoun fixed and consider the template pairs that disambiguate the antecedent as either the occupation or the participant (sentences (a) and (b) in Figure 2). Chance is therefore $0.5^{2}$, or $0.25$.
LingMess has very low pronoun consistency as it struggles with the neopronoun (see Appendix G), and the three smallest FLAN-T5 variants fail at disambiguation consistency, as they tend to mostly predict just the occupation or just the participant. SpanBERT achieves the highest pronoun consistency score (0.60). We take these results to indicate that a) even the best systems are non-robust to pronoun variation, and b) size matters for unsupervised coreference resolution by language models.
5 Measuring Pronominal Bias
Although we have focused on coreference resolution performance and robustness thus far, Winogender schemas were originally proposed to measure gender bias in coreference resolution. They use pronouns, a form of lexical gender, as a proxy for social gender, and correlate incorrect English pronoun resolution with occupational statistics from the USA. However, lexical and social gender are distinct and there are more third person singular pronouns of interest than he and she, and just as coreference resolution systems do not perform as well with pronouns in different grammatical cases, they may also reflect different bias behaviours.
We thus propose a new method for evaluating pronominal bias in coreference resolution with Winogender schemas, correcting for these issues: first, each pair of sentence templates for a given occupation and participant is instantiated with every possible set of pronouns. If the model can correctly disambiguate a pair of templates with a single pronoun like she, then we deem it capable of performing coreference resolution for these templates. If the same model performs differently with the same template when it is instantiated with a different pronoun, e.g., he, there is likely to be bias. If the model over-resolves he to the occupation, we claim that the model has a positive bias between he and that occupation. On the other hand, if it over-resolves he to the participant, we say that the model is biased against associating he with the occupation, i.e., it has a negative bias. In all other cases, we do not have enough information to draw conclusions about bias. This method can be applied to an arbitrary set of pronouns of interest.
We apply our method to SpanBERT to evaluate bias with he/him/his, she/her/her, they/them/their, and xe/xem/xyr. Table 3 shows results for SpanBERT-large, where we find that pronominal bias does not always match across grammatical case, i.e., positive bias with xe for an occupation does not entail positive bias with xyr. In fact, even contradictory associations are possible; SpanBERT-base has a positive bias between manager and them, but a negative bias betweeen manager and their. Across pronoun series, we find that pronominal bias is non-exclusive, i.e., SpanBERT-large has positive bias between baker and her, their as well as xyr, and additionally, these results do not imply that the model must have a negative bias between baker and his. This further highlights the need for evaluation that goes beyond binary, oppositional operationalizations of gender via pronouns.
6 Conclusion
We demonstrate a number of issues with the well-known Winogender schemas dataset, which we fix in our new, expanded WinoPron dataset. In addition, we propose a novel way to evaluate pronominal bias in coreference resolution that goes beyond the binary and focuses on lexical gender. With our new dataset and method, we evaluate both supervised coreference resolution systems and language models, and find that our changes impact performance and bias characteristics. Our work demonstrates that measurements of bias and robustness are only as good as the datasets and metrics we use to measure them, and we call for careful attention when developing resources for and discussing performance on coreference resolution.
Limitations
As in Winogender schemas, our schemas are not “Google-proof” and could conceivably be solved with heuristics, including word co-occurrences, which is a primary concern when creating and evaluating Winograd schemas (Levesque et al., 2012; Amsili and Seminck, 2017; Elazar et al., 2021). The fact that we do not control for this means that our dataset gives generous estimates of model performance and bias, but it also means that this dataset is inappropriate to test model “reasoning.” Our dataset construction instead controls for simple system heuristics that are relevant for coreference resolution, such as always picking the first entity in the sentence, or always picking the second. In interpreting system performance, we also echo Rudinger et al.’s (2018) view of Winogender schemas as having “high positive predictive value and low negative predictive value.” In other words, that they may demonstrate evidence of pronominal bias in systems, but not prove its absence.
We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models (Balloccu et al., 2024). However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on WinoPron with models trained on data after August 2024.
Finally, we note that as our evaluation set only contains one set of templates per occupation-participant pair, our results represent a point in the distribution of bias related to that occupation. In the case of large language models in particular, using a small number of templates for templatic evaluation is known to be brittle even to small, meaning-preserving changes to the template (Seshadri et al., 2022; Selvam et al., 2023). Our dataset’s small size is a result of us requiring a tightly controlled and structured dataset to evaluate how coreference resolution varies. Thus, it may differ from realistic examples (which would have other differences that confound bias results). We wish to emphasize that in addition to controlled datasets like ours, realistic evaluation is also necessary for holistically evaluating performance, robustness and bias in coreference resolution.
Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, and are grateful to Rachel Rudinger and Benjamin Van Durme for their comments. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Related work
Besides Rudinger et al. (2018), there are a number of papers that tackle gender bias in coreference resolution, all of which differ from ours. Similar to Winogender schemas, WinoBias (Zhao et al., 2018) proposes Winograd-like schemas that focus on occupations to evaluate gender bias in coreference resolution. However, WinoBias only covers he and she, rather than our coverage of all English pronoun sets by design. In addition, like Winogender, WinoBias also treats pronouns in all grammatical cases the same way. WinoNB schemas (Baumler and Rudinger, 2022) evaluate how coreference resolution systems handle singular they and plural they with similar schemas. Beyond these constructed schemas, there also exist datasets of challenging sentences found “in the wild,” such as BUG (Levy et al., 2021), GAP (Webster et al., 2018), and GICOREF (Cao and Daumé III, 2021). However, as these natural datasets are not carefully constructed like Winograd-like schemas, pronouns cannot be swapped in dataset instances and still be assumed to be grammatical or coherent.
Appendix B List of Occupations
The occupations along with their respective participants in parentheses are listed below in alphabetical order. This list is identical to the occupations and participants in Rudinger et al. (2018), except that we pair examiner with intern rather than victim:
accountant (taxpayer), administrator (undergraduate), advisor (advisee), appraiser (buyer), architect (student), auditor (taxpayer), baker (customer), bartender (customer), broker (client), carpenter (onlooker), cashier (customer), chef (guest), chemist (visitor), clerk (customer), counselor (patient), dietitian (client), dispatcher (bystander), doctor (patient), educator (student), electrician (homeowner), engineer (client), examiner (intern), firefighter (child), hairdresser (client), hygienist (patient), inspector (homeowner), instructor (student), investigator (witness), janitor (child), lawyer (witness), librarian (child), machinist (child), manager (customer), mechanic (customer) nurse (patient), nutritionist (patient), officer (protester), painter (customer), paralegal (client), paramedic (passenger), pathologist (victim), pharmacist (patient), physician (patient), planner (resident), plumber (homeowner), practitioner (patient), programmer (student), psychologist (patient), receptionist (visitor), salesperson (customer), scientist (undergraduate), secretary (visitor), specialist (patient), supervisor (employee), surgeon (child), teacher (student), technician (customer), therapist (teenager), veterinarian (owner), worker (pedestrian)
Appendix C Annotator Demographics
All three annotators (two authors and an additional annotator) are fluent English speakers. The two authors who create and validate templates have linguistic training at the undergraduate level. One author and one annotator have experience with using singular they and neopronouns, while the other author has prior exposure to singular they but not the neopronoun xe.
Appendix D Annotation Instructions
D.1 Task 1 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 2 data columns and 2 task columns of randomized data. The data columns consist of
  • Sentences which you are asked to annotate for grammaticality; and
  • Questions about pronouns in the sentence, which you are asked to answer
Please be precise in your assignments and do not reorder the data. The columns have built-in data validation and we will perform further tests to check for consistent annotation.
D.1.1 Grammaticality
In the “Grammatical?” column, please enter your grammaticality judgments of the sentence, according to Standard English. The annotation options are:
  • grammatical (for fluent, syntactically valid and semantically plausible sentences)
  • ungrammatical (for sentences that have any typos, grammatical issues, or if the sentence describes a situation that don’t make sense, or just sounds weird)
  • not sure (if you are not sure whether it is clearly grammatical or ungrammatical)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    => grammatical
  • The driver said the passenger that he could pay for the ride with cash.
    => ungrammatical (because ‘said’ is intransitive in Standard English)
D.1.2 Questions about pronouns
Every sentence contains a pronoun, and the “Question” column asks whether it refers to a person mentioned in the sentence or not. The annotation options are:
  • yes (if the pronoun refers to the person)
  • no (if the pronoun does not refer to the person)
  • not sure (if you are not sure about whether the pronoun refers to the person)
Examples:
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the driver?
    => no
  • The driver told the passenger that he could pay for the ride with cash.
    Does the pronoun he refer to the passenger?
    => yes
D.2 Task 2 Description
Together with this annotation protocol, you have received a link to a Google Sheet. The sheet contains 1 randomized data column and 1 task column.
Each row in the data column consists of multiple sentences, of which precisely one sentence contains a blank. Your task is to determine the appropriate pronoun to fill in the blank, and enter it in the “Pronoun” column. Here, appropriate means correct in both form and case.
The tasks are designed to be unambiguous, so please provide only one solution and do not reorder the data.
Example:
  • The driver felt unhappy because he did not make enough money. The driver wondered whether ___ should take out a loan.
    => he
Appendix E Prompting
Table 4 shows all 10 prompt templates we use to present our task instances to FLAN-T5. Each template is presented in three variants to the model, where the options are changed:
  1. 1.
    No options
  2. 2.
    The occupation is presented first and the participant second
  3. 3.
    The participant is presented first and the occupation second
ID Template
[…]
Table 4: Prompting templates, where “task” is filled with each dataset instance, “pronoun” is the only third person singular pronoun that appears in that dataset instance, and “options” is a list of two options to choose from–the occupation and the participant.
Appendix F Data Creation and Validation Details
Constructing the schemas was a challenging and time-consuming process as we ran into intermittent issues, primarily relating to ambiguity (§F.1) and verbal constraints (§F.2), which we had to iteratively fix. In addition to these types of errors, which could only be detected and fixed by humans, we also had structural constraints that could be programmatically checked. We wrote checks for these constraints in Python and describe them in §F.3.
F.1 Ambiguity
One of the big difficulties we faced in creating our schemas was ensuring that we created unambiguous instances of singular specific they. For example, if an advisor and student were meeting to discuss their future, this could also potentially refer to their future together. This problem applied to all grammatical cases (nominative, accusative, and possessive). In addition, possessive sentences were potentially ambiguous across all pronoun series; when discussing a doctor and a patient and someone’s diagnosis, this could be the doctor’s diagnosis (i.e., the diagnosis that was made by the doctor), or the patient’s diagnosis (i.e., the diagnosis that the patient received.
F.2 Verbal Constraints
In some cases the requirement that sentences should be structurally identical until the pronoun created difficulties in finding appropriate, logically and semantically correct endings for the two sentences. It was sometimes difficult to create single-entities sentences that were semantically close to the two-entity versions because the original situation only made sense with two entities (e.g. sentences with "X told Y"). A smaller issue was to ensure that the verbs were congruent with all nominative pronouns. Therefore, we had to only used past tense and avoid the use of was/were.
F.3 Automatic Checks
We automatically checked our data for completeness, i.e., that every occupation-participant pair had sentence templates for nominative, accusative, and possessive grammatical cases. We also checked structural constraints, e.g., that a pair of templates must always be identical until the pronoun slot. We also ensured that there was no ambiguity introduced through additional pronouns appearing in the sentence, limited to the set of pronouns we are interested in, i.e., not including it/its pronouns. Using the human evaluation data, we checked that in every template, the pronoun corefers with what we want it to corefer, according to our data label.
Appendix G Additional Results
Performance on neopronouns (xe/xem/xyr) is markedly worse compared to other pronouns for SpanBERT, CAW-coref and LingMess, but not for the larger sizes of FLAN-T5, as shown in Table 5.
Pronouns LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
he/him/his 0.79 0.80 0.63 0.71 0.52 0.64 0.81 0.89 0.90
she/her/her 0.76 0.77 0.62 0.72 0.52 0.66 0.83 0.91 0.90
they/them/their 0.68 0.64 0.61 0.69 0.52 0.61 0.77 0.89 0.88
xe/xem/xyr 0.09 0.39 0.60 0.69 0.51 0.59 0.70 0.88 0.87
Table 5: F1 scores of coreference resolution systems on WinoPron, split by pronoun series. Red, italicized numbers are worse than chance.
Table 6 shows how coreference resolution systems perform on the single-entity templates in our dataset, which are meant to be a very simple baseline. Here, SpanBERT-base emerges as a clear winner, with near-perfect performance across all cases and pronoun sets.
Data LingMess CAW-coref SpanBERT FLAN-T5
base large small base large xl xxl
All 0.73 0.76 0.96 0.88 0.77 0.76 0.81 0.83 0.84
Nominative 0.80 0.82 0.99 0.99 0.78 0.90 0.85 0.93 0.97
Accusative 0.61 0.65 0.87 0.68 0.76 0.70 0.70 0.70 0.67
Possessive 0.77 0.78 1.00 0.97 0.78 0.78 0.85 0.86 0.89
he/him/his 0.93 0.94 0.95 0.86 0.78 0.81 0.87 0.88 0.89
she/her/her 0.91 0.92 0.96 0.89 0.77 0.81 0.88 0.89 0.87
they/them/their 0.75 0.70 0.96 0.89 0.79 0.76 0.84 0.86 0.87
xe/xem/xyr 0.02 0.27 0.95 0.89 0.75 0.66 0.67 0.69 0.75
Table 6: F1 scores of coreference resolution systems on single-entity sentences in WinoPron. In addition to total F1, we report performance split by grammatical case and by pronoun series.